A coil of wire has 25 loops. Each loop has an area of 0.023 m^2 . A magnetic field is perpendicular to the surface of each of loop at all the times of the magnetic field is changed from 0.23T to 0.34T in 0.24 second. What would be the value of induced emf.
Answer #1

∆B = 0.34 - 0.23 = 0.11 T

Flux

Emf

= (25×0.023× 0.11)/ 0.24

= 0.264 Volt